use Cramer's rule to solve the given system
4x+5y=12
3x+y=-2

Answer by stanbon(57967)

4x+5y=12
3x+y=-2
----------------
Coefficient Determinant = 4-15=-11
x-Determinant =
12 5
-2 1
=12+10=22
So x=22/-11 = -2
----------------
y-Determinant =
4 12
3 -2
=-8-36=-44
So y=-44/-11= 4
-------------------
Comment: When I write this it looks right
but posting may move everything out of place.
Good luck in interpreting what you see.
Cheers,
Stan H.